# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

### v2

@@ 7.1 Parameter identification
Figure 23: Strand magnetization for $\mu_{0}h_{\text{max}}=10$ mT at different frequencies, from virgin state at time $t=0$ to time $t=0.25/f$ (for better clarity). Dashed lines are reference solutions and solid lines are from the ROHM model.
@@ 7.2 Results and numerical performance
Figure 25: Energy loss per cycle for three different field amplitudes as a function of frequency. Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions. The legend is the same for the three subfigures.
Figure 26: Strand magnetization for $\mu_{0}h_{\text{max}}=1$ T at different frequencies. Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions.
@@
8 Conclusion
In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous power loss in composite superconductors, subject to transient transverse external fields. This model is designed by adapting the state-of-the-art energy-based model developed in the context of ferromagnetic hysteresis [27, 28].
We focused specifically on the case of a multifilamentary superconducting strand and proposed two models: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects. We proposed a parameter identification approach for both models and described their inclusion in a finite element framework. Finally, we demonstrated that these models have the potential to reduce the computational cost compared to conventional simulations by several orders of magnitude. The model accuracy can be tailored for specific applications. A very good accuracy (typically with a few percent error on the power loss) in a wide range of field amplitudes and rates of field change can be obtained if required.
This work constitutes a first step towards the homogenization of superconducting magnets, in which the large number of turns makes detailed simulations unrealistically expensive. Replacing the coil windings by homogenized material properties such as the proposed ROHM model allows to describe magnetization and loss efficiently. Further work is necessary to include non-zero transport current, which is not covered in this paper.
Extension of the method to anisotropic systems such as high-temperature superconducting (HTS) tapes, stacks of HTS tapes, or HTS bulks in view of their homogenization can also be considered as further work.
@@ Appendix A Magnetization in 2D Problems
In this section, we justify the introduction of a factor $2$ in the magnetization Eq. (1), in the case of an infinitely long problem solved in 2D with perpendicular currents.
@@
with $\boldsymbol{x}$ the position vector. Consider a rectangular current tube of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig. 28. Accounting for the four sides of the rectangle, the contribution of this current loop to the total magnetic dipole moment is $wl\,\text{d}I$. The two long sides of length $l$ contribute to half of it: $wl\,\text{d}I/2$. The two short sides of length $w$, closing the current loop, contribute to exactly the other half of it: $wl\,\text{d}I/2$.
Figure 28: Current carrying rectangular tube in a cylinder.
When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D, per unit length, on a cross section of the conducting cylinder. To compute the magnetic dipole moment in such a case, however, one must still account for the fact that all current loops must close at the end of the cylinder. These closing currents are not part of the 2D solution, and are then missing from any calculation performed solely on the 2D cross section. Because the closing currents conveniently contribute equally to the magnetic dipole moment compared to the other currents, one can account for them by introducing a factor $2$, as in Eq. (1).
@@ C.1 S cell
For a S cell, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated with Eq. (4.1). The derivative of this update rule is given by
@@ C.2 CS cell
$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (53)
where $\boldsymbol{\mathcal{W}}_{\kappa}$ is the second-order tensor defined in Eq. (C.1), written here as a function of $\boldsymbol{g}_{\text{(p)}}$ instead of $\boldsymbol{h}_{\text{rev(p)}}$.
@@
$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\boldsymbol{\mathcal{W}}_{\chi}(\boldsymbol{g},\boldsymbol{h}_{\text{rev(p)}})\cdot\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (55)

### v3

@@ 7.1 Parameter identification
Figure 23: Strand magnetization for $\mu_{0}h_{\text{max}}=10$ mT at different frequencies, from virgin state at time $t=0$ to time $t=0.25/f$ (for better clarity). Dashed curves are reference solutions. Solid curves are from the ROHM model.
@@ 7.2 Results and numerical performance
Figure 25: Energy loss per cycle for three different field amplitudes as a function of frequency. Solid curves are results from the ROHM model with parameters of Table 1. Dashed curves are reference solutions. The legend is the same for the three subfigures.
Figure 26: Strand magnetization for $\mu_{0}h_{\text{max}}=1$ T at different frequencies. Solid curves are results from the ROHM model with parameters of Table 1. Dashed curves are reference solutions.
@@
8 Conclusion and outlook
In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous power loss induced in composite superconductors by transient transverse magnetic fields. This model is designed by adapting the state-of-the-art energy-based model for ferromagnetic hysteresis [34, 35].
We focused specifically on the case of a multifilamentary superconducting strand and proposed two variations of the ROHM model: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects. We proposed a parameter identification approach for both variations of the model, based on reference solutions obtained with detailed strand models, and we described their inclusion in a finite element framework. Finally, we demonstrated that the ROHM model has the potential to reduce the computational cost compared to conventional simulations by several orders of magnitude. The model accuracy can be tailored for specific applications: a very good accuracy (typically with a few percent error on the power loss) in a wide range of field amplitudes and rates of change can be obtained if required. It also allows to represent deformed strands or strands with defects, provided that reference solutions are used in order to fix adapted ROHM model parameters.
This work constitutes a first step towards the homogenization of superconducting magnet windings, in view of enabling the simulation of magneto-thermal transients in superconducting magnet cross sections [50], accounting for all loss contributions. Replacing the cables with homogenized materials allows to describe the magnet response accurately and efficiently without computing the detailed current density distribution.
The ROHM model accounts for magnetization and loss induced by an external transverse field. One next step consists in defining a second reduced order model to describe the inductance and power loss induced by transport current. The response of a composite strand to transport current is also hysteretic and rate-dependent, therefore, this second reduced order model can be built using similar concepts as those of the ROHM model. Additionally, the ROHM model and the reduced order model for transport current must be coupled. Indeed, the presence of a transport current affects the magnetization, and vice-versa, the presence of a background magnetic field affects the transport current response. Temperature dependence must also be included in the model parameters for magneto-thermal simulations.
For low-temperature superconducting magnets, interstrand coupling current effects in Rutherford cables must also be accounted for. They can be studied using detailed cable models such as in [51] combined with homogenized strand models. The homogenization of these effects into reduced order models will be considered as another next step.
The extension of the approach to high-temperature superconducting (HTS) tapes, stacks of HTS tapes, or HTS bulks in view of their homogenization requires to account for the strong anisotropy of these systems. In principle, as the ROHM model equations are all vectorial, this can be done via the definition of direction-dependent irreversibility parameters and time constants. Further study is required in that direction.
@@ Appendix A Magnetization in 2D Problems
In this section, we justify the introduction of a factor $2$ in the magnetization Eq. (1), in the case of an infinitely long problem solved in 2D with axial currents.
@@
with $\boldsymbol{x}$ the position vector. Consider a rectangular current loop of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig. 28. Accounting for the four sides of the rectangle, the contribution of this current loop to the total magnetic dipole moment is $wl\,\text{d}I$. The two long sides of length $l$ contribute to half of it: $wl\,\text{d}I/2$. The two short sides of length $w$, closing the current loop, contribute to exactly the other half of it: $wl\,\text{d}I/2$.
Figure 28: Rectangular current loop in a cylinder.
When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D, per unit length, on a cross section of the conducting cylinder. To compute the magnetic dipole moment in such a case, however, one must still account for the fact that all current loops must close at the end of the cylinder. These closing currents are not part of the 2D solution, and are therefore missing from any calculation performed solely on the 2D cross section. Because the closing currents conveniently contribute equally to the magnetic dipole moment compared to the other currents, one can account for them by simply introducing a factor $2$, as in Eq. (1).
@@ C.1 S cell
For a S cell, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated with Eq. (27). The derivative of this update rule is given by
@@ C.2 CS cell
$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\tilde{\boldsymbol{\mathcal{W}}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}},\boldsymbol{g}_{\text{(p)}}),$ (52)
where $\tilde{\boldsymbol{\mathcal{W}}}_{\kappa}$ is a second-order tensor similar to that defined in Eq. (C.1), but with $\|\boldsymbol{h}-\boldsymbol{g}_{\text{(p)}}\|\leq\kappa$ as an additional necessary condition for the first case, as in Eq. (28).
@@
$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\boldsymbol{\mathcal{W}}_{\chi}(\boldsymbol{g},\boldsymbol{h}_{\text{rev(p)}})\cdot\tilde{\boldsymbol{\mathcal{W}}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}},\boldsymbol{g}_{\text{(p)}}),$ (54)